# SEAL: Searching Expandable Architectures for Incremental Learning

### v1

@@ 1 Introduction
Incremental Learning (IL), or Continual Learning, is a special set of ML where the models are set to learn a sequential stream of tasks [33]. Under this setting, each task defines a set of observations from which the model needs to learn to perform an action. As the training on new tasks progresses, Deep Neural Networks (DNNs) usually present two unwanted properties: catastrophic forgetting [32] that is the phenomenon of forgetting knowledge from previously seen tasks, and plasticity loss or knowledge saturation [7] that is the inability to keep learning additional tasks. For this reason, many IL solutions focus on mitigating these behaviors. From the stability perspective, some works penalize changes on critical parameters on prior tasks [18, 37], or using knowledge distillation [21]. From the plasticity perspective, promising methods are dynamically expanding networks, that are able to add new weights to the network in order to increase the capacity towards new incoming tasks [36, 27]. IL is particularly useful in many relevant scenarios such as Internet-of-Things and TinyML where the focus is on ensuring efficient adaptation in dynamic and resource-constrained environments [28].
Furthermore, additional work has been developed around the loss landscape geometry, guiding the optimization algorithm towards flatter loss regions to increase robustness of the models, being more resistant to perturbations [9, 29, 15]. These flat solutions have been shown to largely improve the performance in continual learning systems since new incoming tasks can be seen as perturbations of the models [25].
Another direction in ML research is Automated Machine Learning (AutoML), the development of automatic approaches to build more efficient ML pipelines. Within this field, Neural Architecture Search (NAS) aims at automating the design of NN architectures, optimizing them for a specific problem [35]. Many works about NAS have highlighted the importance of flatness to boost the generalization of the searched models [11, 34, 14, 26]. NAS has successfully been applied to IL problems, specifically to automatically search for the optimal model expansion on new tasks arrivals [31, 12, 20]. However, the current approaches perform the search on every task arrival, requiring high computational resources.
This paper proposes an alternative usage for NAS in IL, being the first work jointly searching for the optimal expansion policy and NN architecture. This allows our method to be highly efficient in production, reducing the resources and time needed to train on new task arrivals. Moreover, the method is framed under a specific type of incremental learning problem: a data-incremental scenario [32]. In this context, tasks are defined as disjoint sets of samples, maintaining the same number of classes. We refer to this sets as sample sets, to differentiate the term from traditional IL. The proposed framework, called Searching Expandable Architectures for Incremental Learning (SEAL), utilizes a similar design to FlatNAS [11], using an OFA supernet for faster evaluation of the models [3] and adaptative switching for surrogate prediction [24]. Additionally, in order to accelerate the evaluation and improve the evaluation performance of the expanded models, we define the expansion operator to allow all expanded models to be valid OFA sub-network and perform their evaluation by fine-tuning after initializing the new parts of the expanded models with the pre-trained weights of the OFA supernet.
We tested SEAL on three different datasets and showed that the models retrieved by the framework are competitive with traditional IL methods, while requiring fewer memory resources. Furthermore, the inclusion of a flatness metric showed promising improvements in the IL metrics.
@@
The work is structured as follows. Section 2 introduces the related works regarding NAS solutions for incremental learning. Section 3 provides the background. Section 4 introduces the SEAL framework developed in this study, while in Section 5 the experimental results aiming at evaluating the effectiveness of SEAL are shown. Conclusions are finally drawn in Section 6. To facilitate comparisons and reproducibility, the source code of SEAL is released to the scientific community as a public repository.  [^1] https://github.com/AI-Tech-Research-Lab/SEAL
@@ 2 Related Literature
In this section, we first present techniques for IL and then discuss NAS methods addressing flatness and the incremental problem.
Regarding model expansion, Dynamic Expandable Networks (DEN) [36] were among the first methods to effectively address the plasticity-stability trade-off. DEN enforces sparse weight representations during training to isolate the most relevant parameters for each task. The decision to expand is determined by applying a hard threshold on the pre-computed loss over incoming samples. AdaXpert [27] performs dynamic expansions based on task similarity. The decision to expand is determined using the Wasserstein distance between the distributions of new and previously learned data. The extent of the required expansion is computed from the performance gap. Elastic Weight Consolidation (EWC) [18] mitigates catastrophic forgetting by applying a regularization term to penalize changes to parameters critical for prior tasks, using the Fisher Information Matrix. Synaptic Intelligence (SI) [37] employs a similar regularization strategy but dynamically tracks each weight’s contribution to the loss during training. Learning without Forgetting (LwF) [21] preserves knowledge from previous tasks without requiring old data, using knowledge distillation to retain prior knowledge while learning new tasks. Starting from a static tiny model, Tybox [28] incrementally designs on-device versions of TinyML models while adhering to RAM constraints of target devices. Tybox’s approach seamlessly balances expansion and resource efficiency in memory-constrained environments.
FlatNAS [11], built on a constrained NAS framework [10], introduces a robustness metric grounded in a rescaling-invariant notion of flatness [15, 29] to design architectures resilient to out-of-distribution (OOD) samples while maintaining a constraint on the number of parameters. By incorporating local energy flatness into its objective, which balances accuracy and robustness, FlatNAS effectively guides models toward robust solutions. Other NAS methods leveraging flatness, though not explicitly targeting OOD tasks, include ADVRUSH [26], GeNAS [14], and NA-DARTS [34].
ENAS-S [8] is the first NAS-based solution explicitly tackling the incremental problem, focusing on architectures suitable for edge devices. Their genetic search algorithm optimizes for accuracy and stability, where stability is defined as the accuracy degradation caused by one-shot noise injection in each model layer. ENAS-S demonstrates superior performance in class-incremental tasks compared to traditional CNN models. CLEAS [12] is a NAS using a reinforcement learning controller to have neuron-level control on the expansion search and it is especially effective for catastrophic forgetting as it explicitly isolates previous weights before each expansion. Learn-to-Grow [20] uses DARTS during the search, achieving excellent results. However, the current approaches perform the search on every task arrival, requiring high computational resources. Differently, in our work, we consider all the arrivals in a single search for computation saving and focus specifically on data-incremental tasks.
@@ 3 Technical background
Neural Architecture Search (NAS) solutions can be categorized across three key dimensions [35].
The first dimension is the Search Space, which defines the architecture representation. The simplest approach is the entire-structured Search Space, where the architecture is depicted layer-wise, with each node representing a single layer. Inspired by handcrafted architectures with repetitive motifs [13], the cell-based Search Space defines each node as a cell (or block), which represents a group of layers. Variations of this include hierarchical cells or morphism-based identity transformations between layers.
The second dimension is the Search Strategy, which governs how the search space is explored. Common strategies include reinforcement learning [16], gradient optimization [22], and evolutionary algorithms [2, 1]. Among these, genetic algorithms—particularly NSGA-II [24]—are widely used. NSGA-II generates offspring via specific crossover and mutation operations, selecting the next generation based on fitness through non-dominated sorting and crowding distance comparisons.
The third dimension is the Performance Estimation Strategy, which aims to estimate the performance of candidate architectures without fully training each one, a process that would be computationally expensive. A popular strategy is weight sharing, with Once-For-All (OFA) [3] being a leading example. OFA involves training a supernet that encompasses multiple network configurations once before the search, and candidate networks are evaluated by fine-tuning from the pre-trained supernet weights during every NAS search process. Another approach utilizes surrogate models, such as Gaussian Processes and Radial Basis Functions. For example, MSuNAS [24] introduces an adaptive-switching technique that selects the most accurate predictor among four surrogate models based on a correlation metric, Kendall’s Tau [17], at each NAS iteration.
@@ 3.1 Incremental Learning
IL represents a special kind of ML problem, where models are trained on a sequential stream of tasks, instead of single static dataset. Under this setting, each task consist of a set of observations, following the same or different data distributions, from which the model needs to learn to perform an action. The specific combination of observation and output depends on the type of incremental learning we are solving. IL is considered an open problem for deep learning methods as, experimentally, it has been noted that neural networks are prone to loose previously learned knowledge when training on new tasks [32], an effect known as catastrophic forgetting. Furthermore, their ability to keep learning new tasks is also diminished with the increase of training sessions [7].
@@
$w_{t}^{*}=\min_{w}\sum_{i=0}^{t}\mathcal{L}(D^{*}_{i}|w)$
In this work, we address a problem of data-incremental learning where disjointed dataset samples with same labels arrive sequentially and are not retained in memory for the learner to revisit in the future. Just to mention, other relevant types of settings are domain-incremental, class-incremental, and task-incremental learning [32, 33].
Commonly, there are three relevant aspects when evaluating the performance of continual learning models: overall performance on all seen tasks; knowledge-stability or retaining previously learned knowledge and plasticity for future tasks [33]. Normally, stability and plasticity are opposites in formulation, and a trade-off is required to balance what is more relevant for the model.
Accuracy is the most commonly used metric to measure general performance. Incrementally, we evaluate it with the average accuracy and the average incremental accuracy [4], [33]. The first metric computes how good the model is at solving the tasks for a specific timestamp $t$, while the second adds an average by time, informing the overall progress in time. Consider $a_{ki}$ the accuracy on the task $i$ after training on task $k$, notice $k>i$:
Average Accuracy =$\text{Average Accuracy }=AA_{k}=\dfrac{1}{k}\sum^{k}_{i=0}a_{ki}$
Average Incremental $\text{Average Incremental Accuracy}=AIA_{k}=\dfrac{1}{k}\sum^{k}_{i=0}AA_{k}$
Forgetting is the tendency of the model to replace previously learned tasks with new information. There are two commonly used metrics related to this property: the average forgetting [4] and backward-transfer (BWT) [23], the second measures the influence that learning a new task has for all previously learned tasks.
$\text{Forgetting}=f^{k}_{i}=(\max_{j\in\{0...k-1\}}a_{ji})-a_{ki}$
Average $\text{Average Forgetting}=F_{k}=\dfrac{1}{k}\sum^{k}_{i=0}f^{k}_{i}$
Backward $\text{Backward Transfer}=BWT_{k}=\dfrac{1}{k}\sum^{k}_{i=0}(a_{ki}-a_{ii})$
@@
For a data-incremental setup, the tasks are intrinsically related, so the $\max$ accuracy for a task will not be necessarily immediately after its training session, making average forgetting a more reliable measurement.
Finally, the plasticity of a network relates to its ability to learn new tasks. Although there is no standard measurement for this metric, a common proxy is the forward transfer (FWT) [23] or the influence that learning a task has in future tasks.
Forward $\text{Forward Transfer}=FWT_{k}=\dfrac{1}{k}\sum^{k}_{i=1}(a_{(i-1)i}-\bar{b}_{i})$
Where $\bar{b}_{i}$ represent the accuracy achieved by a randomly initialized model trained on the task $i$. In other words, this metric measures the added performance improvement by the incremental training before the task.
@@ 4.3 Incremental Setting
The measure of the capacity of the current weights to learn the incoming samples is crucial to know when to expand. As in previous works, we measure the cross-entropy loss of the model weights with respect to the incoming samples [36], but instead of applying a threshold on this value, the loss is compared with the previous training iterations. This metric is more robust than setting a threshold directly on the loss value, as the comparison with the previous losses grounds the estimator to a problem the model has already seen and trained on. Formally, given $W_{t-1}$, the weights of the models trained until the last observed sample set and $\mathcal{L}_{t-1}$, the losses stored when receiving the previous sample set. The model is expanded if
$1-\dfrac{\mathcal{L}(W_{t-1},\mathcal{D}_{t})}{\mathcal{L}_{t-1}}<\tau$
@@ 4.4 Expansion of a candidate NN
In an OFA setup, we aim at utilizing as much as possible the weights and architecture already trained on the OFA set. To this end, the constrained OFA set $\Omega_{\bar{x}}$ ensures that each model expansion always produces another model inside the OFA set. This approach reuses OFA weights for expanded sections, reducing fine-tuning time. The weights of the original sections are inherited from the original model. We can specify the expansion of a NN architecture $\bar{x}$ by sampling a 3-elements vector $\vec{d}$ from the IL set $\Omega_{\vec{d}}$. These three bits define three directions on which a model may be expanded, being the depth, width, and kernel size of the candidate NN. Each expansion is applied to the first available OFA-compliant block (e.g., depth expansion can be applied only if the depth of the block is less than the maximum allowed in the OFA set) starting from the last block of layers. The Fig. 2 depicts all the possible expansions in the case of a convolutional layer. In the experimental section, we discuss why the procedure start from the last block instead of the first ones as shown in Fig. 5.
@@ 4.5 Training with cross-distillation
Expanding the model and so increasing model plasticity to learn new samples, nevertheless, may introduce interference in previously learned representations. To avoid this issue, after each expansion, the model is trained with a cross-distillation loss. This forces the added weights to keep a functional representation close to the one before the expansion, reducing catastrophic forgetting [21]. The distillation loss uses the Kullback-Leiber Divergence Loss (KL) on soft targets created by passing the new samples through the old model. The final loss for the expansion follows:
$\mathcal{L}_{CD}=(1-\alpha)\mathcal{L}_{CE}(W_{t},\mathcal{D}_{t})+\alpha\mathcal{L}_{KL}(W_{t-1},D_{t})$
As we are already using the distillation loss to avoid forgetting, the proposed method allows re-training of all model weights after each expansion, thus avoiding the possible problems of keeping frozen the parameters computed on the previous task.

### v2

@@ 1 Introduction
Incremental Learning (IL), or Continual Learning, enables models to learn from a sequential stream of tasks (Wang et al. 2024). In this setting, each task introduces new observations from which the model must learn to perform actions. However, training Deep Neural Networks (DNNs) in this context typically leads to two major challenges: catastrophic forgetting—the tendency to overwrite knowledge from previous tasks (van de Ven, Tuytelaars, and Tolias 2022), and plasticity loss—the model’s inability to learn new information effectively (Dohare et al. 2024). IL methods aim to mitigate these issues by promoting stability (e.g., via regularization (Kirkpatrick et al. 2016; Zenke, Poole, and Ganguli 2017) or knowledge distillation (Li and Hoiem 2017)) and maintaining plasticity (e.g., through dynamic network expansion (Yoone et al. 2017; Niu et al. 2021)). These techniques are especially valuable in resource-constrained scenarios such as IoT and TinyML (Pavone, Lacava, and Brunelli 2024).
Another promising line of work focuses on the geometry of the loss landscape. By guiding optimization toward flatter regions, models become more robust to perturbations (Foret et al. 2021; Pittorino et al. 2021; Jiang et al. 2019), which is beneficial in IL since new tasks can be viewed as perturbations to previously learned solutions (Lyle et al. 2023).
In parallel, Automated Machine Learning (AutoML)—particularly Neural Architecture Search (NAS)—aims to automate the design of neural networks (Wistuba, Rawat, and Pedapati 2019). Flatness has emerged as a desirable property in NAS to enhance generalization (Gambella, Pittorino, and Roveri 2024; Wang et al. 2021; Jeong et al. 2023; Gambella, Pittorino, and Roveri 2025). NAS has also been applied to IL to adapt architectures across tasks (Shahawy, Benkhelifa, and White 2024; Gao et al. 2023; Li et al. 2019), often by searching for optimal expansions. However, most existing approaches repeat the search process at every task, incurring high computational costs.
In this work, we introduce SEAL (Searching Expandable Architectures for Incremental Learning), a novel NAS framework for IL. Unlike prior methods, SEAL jointly searches both the optimal architecture and its expansion policy within a single search phase. This unified approach significantly reduces computational overhead and memory usage while maintaining adaptability across tasks. As a result, SEAL offers a more practical solution for real-world deployment and aligns with key desiderata of NAS for IL—such as efficiency, scalability, and adaptability (Shahawy, Benkhelifa, and White 2025). SEAL specifically targets data-incremental learning, where tasks correspond to disjoint subsets of data while maintaining the same label space (van de Ven, Tuytelaars, and Tolias 2022).
SEAL utilizes a similar design to FlatNAS (Gambella, Pittorino, and Roveri 2024), using an OFA supernet for faster evaluation of the models (Cai et al. 2020) and adaptive switching for surrogate prediction (Lu et al. 2020). To accelerate and improve performance estimation, we constrain the expansion operator so that all expanded architectures remain valid OFA subnetworks. This allows fast evaluation via fine-tuning, using pre-trained OFA weights to initialize newly added components.
We evaluate SEAL on three datasets, showing that it delivers competitive performance with significantly reduced memory requirements compared to traditional IL methods. Moreover, incorporating a flatness-based robustness metric yields consistent improvements in IL metrics.
@@
The remainder of this paper is structured as follows. Section 2 shows the related literature. Section 3 introduces relevant background concepts. Section 4 details the proposed SEAL framework. Section 5 presents experimental results, and Section 6 concludes the paper. To facilitate comparisons and reproducibility, the source code of SEAL is released to the scientific community as a public repository. [^1] https://github.com/AI-Tech-Research-Lab/SEAL
@@ 2 Related Literature
Several techniques have been proposed to address the challenges of incremental learning and architecture robustness. Among the earliest methods tackling model expansion, Dynamic Expandable Networks (DEN) (Yoone et al. 2017) balance plasticity and stability by enforcing sparse weight usage and selectively expanding the network when a loss threshold is exceeded. AdaXpert (Niu et al. 2021) refines this idea by using the Wasserstein distance to measure task similarity and determine when and how much to expand, based on the observed performance gap. Regularization-based methods such as Elastic Weight Consolidation (EWC) (Kirkpatrick et al. 2016) and Synaptic Intelligence (SI) (Zenke, Poole, and Ganguli 2017) mitigate forgetting by penalizing changes to parameters deemed important for previous tasks. Learning without Forgetting (LwF) (Li and Hoiem 2017) takes a different approach by using knowledge distillation to retain prior knowledge without revisiting old data. More recently, Tybox (Pavone, Lacava, and Brunelli 2024) focuses on incrementally adapting TinyML models to meet on-device memory constraints.
In parallel, neural architecture search (NAS) methods have started to incorporate robustness into their objectives. FlatNAS (Gambella, Pittorino, and Roveri 2024) introduces a flatness-based metric inspired by rescaling-invariant formulations (Jiang et al. 2019; Pittorino et al. 2021), guiding the search toward architectures that are robust to distribution shifts. Related works such as ADVRUSH (Mok et al. 2021), GeNAS (Jeong et al. 2023), $A^{2}M$ (Gambella, Pittorino, and Roveri 2025), and NA-DARTS (Wang et al. 2021) further explore flatness as a proxy for generalization and resilience, using strategies such as eigenvalue regularization, random noise injection, or architectural diversity.
A smaller set of NAS methods directly address the incremental learning problem. ENAS-S (Du et al. 2021) applies a genetic search that jointly optimizes for accuracy and stability, defined as robustness to perturbations. CLEAS (Gao et al. 2023) introduces neuron-level expansion through a reinforcement learning controller that isolates old weights before expanding the model. Learn-to-Grow (Li et al. 2019) leverages DARTS to adapt architectures over time and achieves competitive results. ArchCraft (Lu et al. 2024) is a framework that integrates evolutionary algorithms with IL. This approach focuses on optimizing network architectures continuously while preserving previously acquired knowledge. Lastly, PCL  (Lu et al. 2025) proposes a NAS method to design a population of task-specific models rather than a single robust model for IL.
A recent review (Shahawy, Benkhelifa, and White 2025) highlights several pitfalls in current NAS-based IL approaches—namely, redundant retraining, unchecked model growth, and overreliance on replay buffers. SEAL addresses these challenges directly: it avoids retriggering NAS from scratch for each task, expands the architecture only when needed, and removes the dependence on explicit data storage. Furthermore, SEAL exhibits key characteristics desired in IL NAS methods: it supports few-shot learning through cross-task knowledge transfer, maintains a dynamic yet bounded architecture, and introduces mechanisms for graceful forgetting, enabling robustness and scalability across long task sequences.
@@ 3 Technical background
3.1 Neural Architecture Search
Neural Architecture Search (NAS) methods can be categorized along three key dimensions (Wistuba, Rawat, and Pedapati 2019).
The search space defines how architectures are represented. Entire-structured spaces model networks layer by layer, while cell-based spaces, inspired by hand-crafted designs like MobileNets (Howard 2017), group layers into reusable blocks or cells. Extensions include hierarchical structures and morphism-based transformations.
The search strategy determines how the space is explored. Common approaches include reinforcement learning (Kaelbling, Littman, and Moore 1996), gradient-based methods (Liu, Simonyan, and Yang 2019; Xie et al. 2024, 2025), diffusion models (An et al. 2024), Large Language Models (Zeinaty et al. 2025), and evolutionary algorithms (Gambella and Roveri 2023; Gambella et al. 2025). Genetic strategies, especially NSGA-II (Lu et al. 2020; Gambella, Falcetta, and Roveri 2022), are widely adopted for multi-objective optimization.
The performance estimation strategy aims at evaluating NNs without full training, which is computationally infeasible. A common approach is weight sharing, exemplified by Once-For-All (OFA)(Cai et al. 2020), which trains a supernet once and evaluates candidates by fine-tuning its weights. Another strategy uses surrogate models like Gaussian Processes or Radial Basis Functions. For example, MSuNAS(Lu et al. 2020) employs adaptive-switching to select the most accurate predictor among four surrogates using Kendall’s Tau correlation in each NAS iteration.
@@ 3.2 Incremental Learning
Incremental Learning (IL) is a machine learning paradigm where models are trained on a sequence of tasks rather than a fixed dataset. Each task comprises a set of observations—possibly from different distributions—used to learn specific actions, depending on the IL setting. IL poses a challenge for deep learning models, which tend to forget previous tasks when learning new ones—a phenomenon known as catastrophic forgetting (van de Ven, Tuytelaars, and Tolias 2022). Additionally, their capacity to learn new tasks declines as more training sessions are performed (Dohare et al. 2024).
@@
$w_{t}^{*}=\min_{w}\sum_{i=0}^{t}\mathcal{L}(D^{*}_{i}|w)$ (1)
In this work, we address a problem of data-incremental learning where disjointed dataset samples with same labels arrive sequentially and are not retained in memory for the learner to revisit in the future. Just to mention, other relevant types of settings are domain-incremental, class-incremental, and task-incremental learning (van de Ven, Tuytelaars, and Tolias 2022; Wang et al. 2024).
IL performance is typically evaluated across three key aspects: overall performance on all tasks, stability (retention of past knowledge), and plasticity (adaptability to new tasks) (Wang et al. 2024).
Accuracy remains the standard metric. In IL, we report both average accuracy and average incremental accuracy (Chaudhry et al. 2018; Wang et al. 2024). The former reflects task performance at a given time, while the latter summarizes progress across all tasks. Let $a_{ki}$ be the accuracy on task $i$ after learning task $k$ ($k\geq i$):
$\text{Average Accuracy}\quad AA_{k}$ $=\frac{1}{k}\sum_{i=0}^{k}a_{ki}$ (2)
$\text{Average Incremental Accuracy}\quad AIA_{k}$ $=\frac{1}{k}\sum_{i=0}^{k}AA_{k}$ (3)
Forgetting quantifies knowledge loss, while backward transfer (BWT) measures the impact of new learning on previous tasks (Chaudhry et al. 2018; Lopez-Paz and Ranzato 2017):
$\text{Forgetting}\quad f^{k}_{i}$ $=\left(\max_{j\in\{0,\dots,k-1\}}a_{ji}\right)-a_{ki}$ (4)
$\text{Average Forgetting}\quad F_{k}$ $=\frac{1}{k}\sum_{i=0}^{k}f^{k}_{i}$ (5)
$\text{Backward Transfer}\quad BWT_{k}$ $=\frac{1}{k}\sum_{i=0}^{k}\left(a_{ki}-a_{ii}\right)$ (6)
@@
In data-incremental settings—where tasks are highly related—the max accuracy for a task may not occur immediately after learning it, making average forgetting a more robust metric.
Forward transfer (FWT) approximates plasticity by measuring the influence of past learning on future tasks (Lopez-Paz and Ranzato 2017):
$FWT_{k}=\dfrac{1}{k}\sum^{k}_{i=1}(a_{(i-1)i}-\bar{b}_{i})$ (7)
where $\bar{b}_{i}$ is the accuracy of a randomly initialized model trained only on task $i$. FWT thus reflects the performance gain due to prior experience.
@@ 4.3 Incremental Setting
To decide when to expand the model, we assess its capacity to learn new data by comparing the current cross-entropy loss to that of previous training iterations. Unlike fixed thresholds used in prior works (Yoone et al. 2017), this relative comparison provides a more robust and adaptive criterion, as it reflects how well the model handles newly arrived data based on past performance. Formally, given the weights $W_{t-1}$ and loss $L_{t-1}$ from the previous sample set, the model is expanded if
$1-\dfrac{\mathcal{L}(W_{t-1},\mathcal{D}_{t})}{\mathcal{L}_{t-1}}<\tau$ (10)
@@ 4.4 Expansion of a candidate NN
In an OFA setup, we reuse the weights and structure of the pre-trained OFA supernet as much as possible. To ensure compatibility, the constrained OFA set $\Omega_{\bar{x}}$ restricts expansions to produce valid sub-networks within the OFA space. Each expansion inherits weights from the original model and applies changes to the first available OFA-compliant block, starting from the last layers. The expansion is defined by a 3-bit vector $\vec{d}$ indicating changes in depth, width, and kernel size. Figure 2 illustrates these operations, while the rationale for expanding from the last block is discussed in Figure 5.
@@ 4.5 Training with cross-distillation
Expanding the model and so increasing model plasticity to learn new samples, nevertheless, may introduce interference in previously learned representations. To avoid this issue, after each expansion, the model is trained with a cross-distillation loss. This forces the added weights to keep a functional representation close to the one before the expansion, reducing catastrophic forgetting (Li and Hoiem 2017; Nogales, Gambella, and Roveri 2025). The distillation loss uses the Kullback-Leiber Divergence Loss (KL) on soft targets created by passing the new samples through the old model. The final loss for the expansion follows:
$\mathcal{L}_{CD}=(1-\alpha)\mathcal{L}_{CE}(W_{t},\mathcal{D}_{t})+\alpha\mathcal{L}_{KL}(W_{t-1},D_{t})$ (11)